Documenta Mathematica

Extra Volume: Kazuya Kato’s Fiftieth Birthday, 2003

Preface  1
Prime Numbers  3

Ahmed Abbes and Takeshi Saito Ramification of Local Fields

with Imperfect Residue Fields II  5–72

Kenichi Bannai Specialization of the p-adic Polylogarithm to

p-th Power Roots of Unity  73–97

Laurent Berger Bloch and Kato’s Exponential Map:

Three Explicit Formulas  99–129

Spencer Bloch and Hélène Esnault

The Additive Dilogarithm  131–155

David Burns and Cornelius Greither Equivariant Weierstrass Preparation

and Values of L-functions at Negative Integers  157–185

John Coates, Peter Schneider, Ramdorai Sujatha Links Between Cyclotomic

and GL_{2} Iwasawa Theory  187–215

Robert F. Coleman

Stable Maps of Curves  217–225

Christopher Deninger Two-Variable Zeta Functions

and Regularized Products  227–259

Ivan Fesenko

Analysis on Arithmetic Schemes. I  261–284

Jean-Marc Fontaine

Presque C_{p}-Représentations  285–385

Takako Fukaya Coleman Power Series for K_{2} and p-Adic Zeta Functions

of Modular Forms  387–442

Y. Hachimori and O. Venjakob Completely Faithful Selmer Groups

over Kummer Extensions  443–478

Uwe Jannsen and Shuji Saito Kato Homology of Arithmetic Schemes and Higher Class Field Theory

over Local Fields  479–538

Masato Kurihara On the Structure of Ideal Class Groups

of CM-Fields  539–563

Nobushige Kurokawa, Hiroyuki Ochiai, and Masato Wakayama

Absolute Derivations and Zeta Functions  565–584

Barry Mazur and Karl Rubin

Studying the Growth of Mordell-Weil  585–607

Shinichi Mochizuki The Absolute Anabelian Geometry

of Canonical Curves  609–640

Hyunsuk Moon and Yuichiro Taguchi Refinement of Tate’s Discriminant Bound and Non-Existence Theorems for Mod p

Galois Representations  641–654

Arthur Ogus On the Logarithmic

Riemann-Hilbert Correspondence  655–724

Bernadette Perrin-Riou Groupes de Selmer et Accouplements; Cas Particulier des

Courbes Elliptiques  725–760

Christophe Soulé A Bound for the Torsion in the K-Theory

of Algebraic Integers  761–788

Akio Tamagawa Unramified Skolem Problems and Unramified Arithmetic Bertini Theorems

in Positive Characteristic  789–831

Takeshi Tsuji On the Maximal Unramified Quotients of p-Adic Étale Cohomology Groups

and Logarithmic Hodge–Witt Sheaves  833–890

Nobuo Tsuzuki On Base Change Theorem

and Coherence in Rigid Cohomology  891–918